OpenStudy (anonymous):

10*9*8*7*6*5*4*3*2*1 = 10! Can this be true?! Why or why not?

OpenStudy (amistre64):

10! is defined as the product of 10 and its preceding whole numbers

OpenStudy (amistre64):

why is it true? becasue thats how we define it lol

OpenStudy (anonymous):

The expression "x!" means x(x-1)(x-2)(x-3)...1 So 10! means 10(10-1)(10-2)(10-3)(10-4)(10-5)(10-6)(10-7)(10-8)(10-9)= 10*9*8*7*6*5*4*3*2*1=10!, which means it is correct.

OpenStudy (amistre64):

the 1 on the end is redundant....factual, but redundant

OpenStudy (anonymous):

Why wouldn't it be true though?